it says find Sn for each arithmetic series described. a1=20, an=112, d=4. how do i find it ?
Answer by stanbon(75887)

it says find Sn for each arithmetic series described. a1=20, an=112, d=4. how do i find it ?
-------------------------
S(n) is the sum of n terms of the series.
The formula is S(n) = (n/2)(a(1)+a(n))
-----------------------------------
You need to find "n"
a(n) = a(1)+(n-1)d
112 = 20 + (n-1)*4
92 = (n-1)4
n-1 = 23
n = 24
---
So you are looking for the sum of the 1st 24 terms of the series.
---
Use the formula to get: S(24) = (24/2)(20+112) = 12*132 = 1584
====================================================================
Cheers,
Stan H.
Reply to stanbon@comcast.net